OpenStudy (anonymous):

is (8,-6) a solution of 8x-6y=7 I got NO

OpenStudy (anonymous):

8(8)-6(-6)=64+36=100 which is not equal to 7 so yeah it is not.

OpenStudy (anonymous):

yea.its not a correct solution ...........

OpenStudy (anonymous):

u can check by jst plug in value =0 for both x,y separately........